Mry has net and taxable income of $55000; her claim for charitable donation is $4000. How much she is saving tax ?
Answer #1

Tax bracket= 4453.5 + 22%*( Taxable income- 38700). This is according to the new tax bracket in US, 2019.

Without donation: Taxable income= $55000

Tax= 4453.5+22%(55000-38700) = $8039.5

With donation under Schedule A Taxable income= $51000

Tax= 4453.5+(51000-38700)*22% = $7159.5

Savings = $8039.5 - $7159.5 = $880